# Flow-Modulated Scoring for Semantic-Aware Knowledge Graph Completion

### v2

@@
Abstract.
Effective modeling of multifaceted relations is pivotal for Knowledge Graph Completion (KGC). However, a majority of existing approaches are predicated on static, embedding-based scoring, exhibiting inherent limitations in capturing contextual dependencies and relational dynamics. Addressing this gap, we propose the Flow-Modulated Scoring (FMS) framework. FMS comprises two principal components: (1) a semantic context learning module that encodes context-sensitive entity representations, and (2) a conditional flow-matching module designed to learn the dynamic transformation from a head to a tail embedding, governed by the aforementioned context. The resultant predictive vector field, representing the context-informed relational path, serves to dynamically refine the initial static score of an entity pair. Through this synergy of context-aware static representations and conditioned dynamic information, FMS facilitates a more profound modeling of relational semantics. Comprehensive evaluations on several standard benchmarks demonstrate that our proposed method surpasses prior state-of-the-art results.
Knowledge graph, Graph embedding, Knowledge graph completion, Diffusion model, Graph representation learning, Graph neural network
ccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Machine learningcopyright: none
1. Introduction
[…]
Figure 1. This figure systematically depicts the dynamic interplay of Iron Man’s relationships with War Machine, Captain America, and the Winter Soldier, all broadly defined as allies. However, the specific semantics of these alliances exhibit varying degrees, which are profoundly shaped by the semantic context—namely, distinct historical junctures, critical incidents, and individual value perspectives.
@@ 1. Introduction
However, due to the continuous evolution of knowledge and the high cost and complexity of knowledge acquisition, real-world knowledge graphs, despite their vast scale, often suffer from incomplete or missing information. This limitation partially undermines their application efficacy and knowledge inference capabilities. To address this challenge, the task of Knowledge Graph Completion (KGC) has emerged, aiming to infer and predict missing links based on existing facts.
Relation Prediction(RP) between entity pairs stands as one of the pivotal sub-tasks within Knowledge Graph Completion. Specifically, given a head entity $\boldsymbol{h}$ and a tail entity $\boldsymbol{t}$, the RP task aims to learn a function $f(\cdot)$ that predicts a set of plausible relations $\mathcal{R}$ between them, formally denoted as $f(\boldsymbol{h},\boldsymbol{t})=\mathcal{R}$. To address this, researchers have explored a multitude of methodologies. Initially, embedding-based models, such as TransE (Bordes et al., 2013), were pioneered. These models conceptualize relations as translation vectors from the head entity to the tail entity, evaluating the plausibility of a triplet via simple vector operations. Despite their conciseness and efficiency, such methods often struggle to effectively handle complex relational patterns, including one-to-many, many-to-one, and many-to-many scenarios. To overcome these limitations, subsequent works such as TransH (Wang et al., 2014), TransR (Lin et al., 2015) and TransD (Ji et al., 2015) significantly improved the modeling capacity for complex relations by introducing relation-specific projection hyperplanes or mapping matrices. Concurrently, another line of research has focused on leveraging more sophisticated semantic matching architectures. For instance, models based on multilinear interactions such as DistMult (Yang et al., 2015), and ComplEx (Trouillon et al., 2016), aim to capture deeper and richer interaction patterns between entities and relations by designing more elaborate scoring functions. While these embedding-based methods have achieved considerable progress in KGC tasks, their core paradigm predominantly relies on learned static embeddings for entities and relations, employing a fixed scoring function to assess the confidence of triplets. This paradigm may, to some extent, fall short in fully capturing the subtle semantic nuances of relations across diverse contexts and the dynamic evolutionary characteristics of inter-entity relationships, thereby potentially limiting their representational power and predictive accuracy. Taking the relationships among Marvel characters as an example (see the Figure 1), in a work-related context, Iron Man and War Machine share a military liaison relationship, with War Machine acting as Iron Man’s supporter, making them core allies. Iron Man and Captain America are co-founders of the Avengers but hold opposing views on the Sokovia Accords; overall, they remain key allies. In contrast, the Winter Soldier, despite being Captain America’s best friend, assassinated Iron Man’s parents, resulting in a complex ally relationship with Iron Man. Clearly, these characters’ relationships vary across semantic contexts, with differing degrees of closeness. Notably, during reasoning, the semantic contexts we rely on are more about relational attributes than entity attributes.
Since then, GNN-based methods have gradually emerged. For instance, in the RP task, the most representative method, PathCon (Wang et al., 2021), models the probability distribution of relations by incorporating the context and multi-hop paths of entity pairs to achieve prediction. However, like other GNN-based methods, this approach is constrained by static context, thus failing to effectively handle the dynamic nature of relations.
Recently, inspired by the strong semantic comprehension of large language models (LLMs), LLM-based approaches have been increasingly applied to relation prediction tasks. For instance, KG-BERT (Yao et al., 2019) converts knowledge graph triples into textual sequences, leveraging BERT’s semantic encoding for relation prediction. Nadkarni et al. (Nadkarni et al., 2021) employ a BERT model fine-tuned for the biomedical domain to effectively capture complex relational patterns among biomedical entities. Similarly, Nassiri et al. (Nassiri et al., 2022) introduce a triplet BERT-network model, enhancing relation prediction accuracy through semantic refinement. While these methods have shown notable advances in discerning subtle semantic nuances of relations across diverse contexts, they fail to explicitly model the dynamic evolution of inter-entity relationships. Furthermore, the large parameter scale of LLMs can hinder inference efficiency, increase deployment costs, and limit adaptability to domain-specific knowledge graphs. To address all these challenges, we propose the Flow-Modulated Scoring (FMS) framework. This framework integrates a semantic context learning module to encode context-aware embeddings and employs a condition flow-matching module to model dynamic transformations between head and tail entities. By generating a context-driven predictive vector field, FMS dynamically adjusts static scores, thereby achieving more comprehensive and precise modeling of relation semantics across diverse contexts. Our contributions can be summarized as follows:
  • We propose the Flow-Modulated Scoring (FMS) framework for knowledge graph completion, which synergizes context-aware embeddings with dynamic relation modeling, offering superior expressiveness and accuracy compared to static embedding-based methods;
  • We develop a semantic context learning module that approximates semantic similarity via an energy function, selectively aggregating the Top-k relation contexts from randomly sampled candidates to construct expressive entity representations;
  • We introduce a condition flow-matching module that learns a dynamic vector field to model the transformation from head to tail entity embeddings, employing Brownian Bridge Matching to capture both deterministic and stochastic relational dynamics;
  • We demonstrate through comprehensive experiments that FMS achieves state-of-the-art performance across all datasets, validating its robustness in capturing dynamic relation semantics and superiority in parameter.
2. Related Works
2.1. Knowledge Graph Completion
Knowledge Graphs provide structured information for downstream tasks like recommender systems and semantic analysis. Knowledge Graph Completion, particularly Relation Prediction (RP), addresses incomplete KGs by predicting missing links. Most KGC methods rely on embedding-based approaches, which assign vectors to entities and relations in a continuous space, trained on observed facts, such as TransE (Bordes et al., 2013), treat entities as points and relations as translations, aiming to position the translated head entity close to the tail entity in real, complex, or quaternion spaces. These methods effectively handle multiple relation patterns and achieve strong results. Alternatively, bilinear models like DistMult (Yang et al., 2015) and ComplEx (Trouillon et al., 2016) compute semantic similarity via matrix or vector dot products. Some methods explore advanced architectures beyond point vectors, such as CNN-based models (Dettmers et al., 2018). However, these embedding-based approaches struggle with inductive settings. Graph Neural Network GNN-based methods, like PathCon (Wang et al., 2021), incorporate relation contexts and multi-hop paths but often neglect semantic consistency and dynamic relation modeling. Our work proposes a dynamic, context-aware framework to overcome these limitations.
Figure 2. The FMS architecture.
2.2. Diffusion Models
Diffusion models, inspired by non-equilibrium thermodynamics, have achieved remarkable success in fields such as computer vision (Li et al., 2022b), sequence modeling (Li et al., 2022a; Tashiro et al., 2021), and audio processing (Chen et al., 2020; Kong et al., 2020). Predominant approaches include Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) and Score-based Generative Models (SGMs) (Song et al., 2020). SGMs utilize Stochastic Differential Equations (SDEs) for generation, and DDPMs can be considered an instance of SGMs under specific discretization schemes. Meanwhile, Continuous Normalizing Flows (CNFs) (Chen et al., 2018) are another attractive class of generative modeling techniques, learning an invertible mapping between the data distribution and a simple prior distribution via neural network-parameterized Ordinary Differential Equations (ODEs). Recently, research has shown that CNFs can also be trained using a regression objective similar to those in diffusion models, a method known as Flow Matching (FM) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2023; Liu, 2022). FM learns the transformation by directly regressing the ODE’s vector field (drift term), which is analogous to how diffusion models train score functions. The advent of FM introduced new ideas for CNF training and improved sample quality, though early FM methods typically assumed a Gaussian source distribution, a limitation addressed in subsequent work (Chen and Lipman, 2024; Pooladian et al., 2023). Against this backdrop, Condition Flow-Matching (CFM) (Tong et al., 2023a) was proposed as a generalization of FM models. CFM aims to provide a unified, simulation-free training framework for FM models with arbitrary transport maps. It inherits the stability of regression objectives from diffusion models while striving for efficient inference of deterministic flow models. Unlike traditional diffusion models and early CNF training algorithms, CFM does not require the source distribution to be Gaussian, nor does it need to evaluate its probability density.
3. Methodology
3.1. Preliminaries
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an instance of a knowledge graph, where $\mathcal{V}$ is the set of entity nodes and $\mathcal{E}$ is the set of edges, with each edge representing a relation $r\in\mathcal{R}$. Our goal is to predict missing relations in the knowledge graph, i.e., given an entity pair $(h,t)$, infer the relation $r$ between them. Specifically, we aim to model the distribution of relation types given the entity pair $(h,t)$, denoted as $p(r|h,t)$, to capture the complex semantic context and dynamic nature of relations. According to Bayes’ theorem, this distribution can be expressed as:
(1) $p(r|h,t)\propto p(h,t|r)\cdot p(r),$
@@
(2) $p(h,t|r)=\frac{1}{2}\left(p(h|r)\cdot p(t|h,r)+p(t|r)\cdot p(h|t,r)\right).$
@@
Equation (2) provides guidance for our modeling approach. The terms $p(h|r)$ and $p(t|r)$ measure the likelihood of entities $h$ or $t$ given the relation $r$, reflecting the semantic context of the entities. The terms $p(t|h,r)$ and $p(h|t,r)$ capture the dynamic transformation process from $h$ to $t$ or from $t$ to $h$ given the relation $r$, reflecting the dynamic nature of relations. Below, we will illustrate how these two factors are modeled within our approach and how they contribute to relation prediction.
3.2. Flow-Modulated Scoring Model
@@
Figure 2 illustrates the architecture of the Flow-Modulated Scoring model. It consists of two stages: a semantic context learning module and a condition flow-matching module. The first stage encodes context-aware entity embeddings, while the second learns dynamic transformations from head to tail embeddings, generating a context-informed vector field. This field modulates the static score of entity pairs, with the ultimate goal of achieving comprehensive relation prediction. Next, we provide a detailed explanation of these modules.
3.2.1. Semantic Context Learning Module

### v3

@@
Abstract
Knowledge graph completion demands effective modeling of multifaceted semantic relationships between entities. Yet, prevailing methods, which rely on static scoring functions over learned embeddings, struggling to simultaneously capture rich semantic context and the dynamic nature of relations. To overcome this limitation, we propose the Flow-Modulated Scoring (FMS) framework, conceptualizing a relation as a dynamic evolutionary process governed by its static semantic environment. FMS operates in two stages: it first learns context-aware entity embeddings via a Semantic Context Learning module, and then models a dynamic flow between them using a Conditional Flow-Matching module. This learned flow dynamically modulates a base static score for the entity pair. By unifying context-rich static representations with a conditioned dynamic flow, FMS achieves a more comprehensive understanding of relational semantics. Extensive experiments demonstrate that FMS establishes a new state of the art across both canonical knowledge graph completion tasks: relation prediction and entity prediction. On the standard relation prediction benchmark FB15k-237, FMS achieves a near-perfect MRR of 99.8% and Hits@1 of 99.7% using a mere 0.35M parameters, while also attaining a 99.9% MRR on WN18RR. Its dominance extends to entity prediction, where it secures a 25.2% relative MRR gain in the transductive setting and substantially outperforms all baselines in challenging inductive settings. By unifying a dynamic flow mechanism with rich static contexts, FMS offers a highly effective and parameter-efficient new paradigm for knowledge graph completion. Code published at: https://github.com/yuanwuyuan9/FMS.
Index Terms:
Knowledge Graph, Flow-Matching, Graph Neural Network, Knowledge Representation and Reasoning.
1 Introduction
@@
Knowledge Graphs (KGs), as a robust formalism for knowledge representation, have become pivotal in a myriad of large-scale applications, including recommender systems [1, 2], information extraction [3], and question answering [4, 5], owing to their remarkable efficacy in organizing and representing complex semantic facts [6, 7, 8, 9]. However, real-world KGs, despite their vast scale, are ubiquitously plagued by incompleteness [10, 11, 12]. This issue stems from the continuous evolution of knowledge itself and the high cost and complexity associated with its acquisition. To address this fundamental challenge, the task of Knowledge Graph Completion (KGC) has emerged, which aims to infer and predict missing links based on existing facts [13, 14, 15, 16].
Figure 1: The context-dependent nature of relational semantics, exemplified by Iron Man’s alliances. The generic relation ”ally” acquires distinct meanings (e.g., Highest Core, High Key, Complex Conditional) based on specific contextual facts, such as shared history or critical conflicts.
Figure 2: Illustration of the dynamic evolution of relationships. The examples show the changing dynamics over time between (left) Thor and Loki, and (right) Iron Man and Spider-Man, highlighting that relationships progress through distinct stages.
Among the sub-tasks of KGC, Relation Prediction (RP) is crucial for enabling fine-grained knowledge reasoning [17, 18]. It is dedicated to inferring the set of potential relations $\mathcal{R}$ for a given entity pair $(h,t)$ [19]. Existing methodologies for this task predominantly fall into two categories. The first category comprises embedding-based models, such as TransE [20], TransH [21], and DistMult [22], which assess the plausibility of a triplet by learning static entity and relation embeddings within a fixed scoring function. The second category consists of GNN-based methods that aggregate contextual information [23, 24], like PathCon [19], which leverages multi-hop neighborhood structures to enrich semantic representations. Despite the progress achieved, these methods are constrained by two fundamental bottlenecks:
1) Inability to Capture Rich Contextual Information Shaping Relation Semantics. The meaning of a relation is often highly contingent on its surrounding factual context [19, 25, 15]. As illustrated in Figure 1, the generalized label “ally” is insufficient to accurately describe the relationships between Iron Man and other characters. His relationship with War Machine, rooted in their military liaison, is a “core, top-level alliance,” whereas his relationship with Captain America is a ”key, high-level alliance” shaped by their shared history of team-building and subsequent political disagreements. This indicates that effective relation prediction must account for the subtle contextual nuances that differentiate superficially similar yet substantially distinct relations. 2) Incapacity of Static Paradigms to Model the Dynamic Evolution of Relations. In reality, relationships exhibit path-dependent and time-evolving characteristics [9, 26]. For instance, the Thor-Loki dynamic shifted from brotherhood to enmity and later to reconciliation, while the Iron Man-Spider-Man bond evolved from a mentor-protégé dynamic to a quasi-paternal one (see Figure 2). Accurately modeling such evolutionary trajectories is critical, a task for which existing static scoring functions are ill-equipped. While GNNs introduce structural context, they operate on fixed snapshots of the graph and are thus inherently static in their modeling approach [19].
Recently, Large Language Models (LLMs) have been introduced to enhance semantic understanding in KGC [27, 10, 28], as seen in models like KG-BERT [29] and MuKDC [30]. Although these models offer improved context awareness, they still lack explicit mechanisms for modeling the evolutionary paths of relations and are often hampered by large parameter counts, low efficiency, and challenges in domain adaptation. Furthermore, approaches like MuKDC are also constrained by their strong reliance on meticulously designed prompts[9, 31].
To this end, this paper introduces the Flow-Modulated Scoring (FMS) framework, a novel approach that pioneers the unification of context-aware and dynamic evolution modeling for knowledge graph completion. The core idea of FMS is to conceptualize a relation as a conditional vector flow from a head entity $h$ to a tail entity $t$, where the trajectory of this flow is dynamically modulated by the local semantic context. Specifically, the FMS framework comprises two key modules: 1) Semantic Context Learning Module employs an energy function to identify and aggregate the top-$k$ most semantically relevant contextual paths, thereby constructing context-aware entity representations. 2) Conditional Flow Matching Module learns a conditional vector field that explicitly models the dynamic evolutionary trajectory of a relation from $h$ to $t$, enabling a path-sensitive representation of relational semantics.
By integrating these components, FMS dynamically adjusts the output of conventional scoring functions, facilitating a fine-grained modeling of relation semantics that is both context-sensitive and temporally aware. Extensive experiments demonstrate that FMS achieves state-of-the-art (SOTA) performance across multiple standard benchmarks while maintaining superior parameter efficiency. The main contributions of this paper are summarized as follows:
  • Flow-Modulated Scoring Framework: We propose the first KGC framework that integrates context-aware and dynamic evolution modeling, overcoming the limitations of static scoring paradigms (Section 3.2).
  • Semantic Context Learning: We design an energy-based top-$k$ context selection mechanism to efficiently capture critical semantic dependencies (Section 3.2.1).
  • Conditional Flow Matching: We introduce a conditional vector flow mechanism to explicitly model the evolutionary paths of relations, enhancing the capacity for dynamic semantic representation (Section 3.2.2).
  • We demonstrate the state-of-the-art performance and high parameter efficiency of FMS on multiple benchmarks, confirming its effectiveness in modeling complex relations (Section 4 and Section 5).
2 Related Work
2.1 Knowledge Graph Completion
Knowledge Graphs provide structured information for downstream tasks such as recommender systems and semantic analysis [13, 9, 32]. Knowledge Graph Completion (KGC), and particularly Relation Prediction (RP), addresses the problem of KG incompleteness by predicting missing links [10, 33, 34, 35].
The majority of KGC methods rely on embedding-based approaches, which assign vectors to entities and relations in a continuous space and are trained on observed facts [36, 37, 38, 39]. For instance, early geometric and bilinear models, such as TransE [20] and ComplEx [40], laid the groundwork by defining simple yet effective scoring functions. To capture more intricate relational patterns, subsequent research has advanced into more expressive geometric spaces. RotaE [14], for example, introduced relational rotations in complex space, a concept later enhanced by Rot-Pro [41] for greater fidelity. Other approaches have explored specialized geometries, such as HAKE’s [42] use of polar coordinates for semantic hierarchies and BoxE’s [43] hyper-rectangle representations for logical conjunctions. In a parallel thread of research, recent work has focused not on geometry but on augmenting the reasoning process itself. AnKGE [44] exemplifies this direction by equipping base models with an analogical inference mechanism to leverage similar, known facts [24].
Despite their effectiveness in handling relational patterns, these models are fundamentally static. They rely on pre-computed, fixed embeddings, which renders the models context-agnostic; the representation of an entity or relation remains constant regardless of the specific factual query [9]. This paradigm is inherently incapable of capturing the subtle nuances and contextual dependencies of relational meanings. To overcome static embeddings, recent research has leveraged Large Language Models (LLMs) to inject contextual awareness [45, 35]. These approaches either refine link prediction as a textual scoring task (e.g., KG-BERT [29], SimKGC [46]) or utilize LLMs as generative knowledge sources to distill missing facts, as exemplified by MuKDC [30]. Despite their promise, these LLM-based methods are often hampered by their massive parameter counts, low computational efficiency, and significant challenges in domain adaptation [35]. Another direction to overcome static embeddings is through generative modeling. Recent works like FDM [37] and KGDM [47] employ diffusion models [48] to learn fact distributions. However, their generative process is typically conditioned solely on the local query triplet, thus failing to leverage the broader, multi-hop structural context crucial for robust link prediction.
The advent of GNN-based models represents an attempt to incorporate contextual information from the graph structure [49, 50]. Methods like PathCon [19] enrich entity and relation representations by aggregating information from multi-hop paths, demonstrating the value of neighborhood context. To mitigate the high cost of explicit path extraction, recent work CBLiP [51] instead uses connection-biased attention to implicitly capture structural information from the local subgraph. However, these approaches still operate on a static snapshot of the graph. While they encode structural context, they do not explicitly model the formation process or evolutionary trajectory of relations. Their aggregation mechanisms are typically fixed and cannot adapt to the dynamic and path-dependent characteristics inherent in how relationships form and evolve. In summary, prevailing methodologies exhibit an over-reliance on static modeling paradigms, making them ill-equipped to handle the dual challenges of context-sensitivity and dynamic evolution.
2.2 Modeling Dynamic Processes with Conditional Flow Matching
To address the core challenges of modeling dynamic relationship evolution, we draw on research in generative models concerning the learning of transformations and trajectories between data distributions. Although diffusion models [48, 52] have demonstrated formidable generative capabilities in domains such as computer vision [53] and sequence modeling [54, 55], their training and inference processes are often computationally expensive.
A more direct and efficient approach for learning transformation paths is offered by Continuous Normalizing Flows (CNFs) [56]. CNFs learn an invertible mapping between a data distribution and a simple prior distribution by parameterizing the vector field of an Ordinary Differential Equation (ODE). Recently, Flow Matching (FM) [57, 58, 59, 60] has emerged as a highly efficient and stable technique for training CNFs. Unlike score-matching methods that learn the gradient of the data distribution’s logarithm, FM learns the transformation by directly regressing the ODE’s vector field. This simulation-free training approach significantly enhances model stability and efficiency. The advent of FM introduced new paradigms for CNF training and improved sample quality, and while early FM methods often assumed a Gaussian source distribution, this limitation has been addressed in subsequent work [61, 62].
Building on this foundation, Conditional Flow Matching (CFM) [63, 64, 65] offers a more generalized and powerful framework. CFM aims to learn probability paths conditioned on specific attributes, providing a unified, simulation-free training objective for models with arbitrary transport maps. Crucially, it neither requires the source distribution to be Gaussian nor necessitates the evaluation of its probability density. These properties—efficiency, stability, and flexible conditioning—make CFM the ideal theoretical tool for our objectives. In this work, we innovatively apply the core mechanism of Conditional Flow Matching within KGC. Instead of using it for data generation, we leverage it to model the conditional evolutionary path of a relation from a head entity $h$ to a tail entity $t$. This enables us to explicitly model relations as trajectories, dynamically modulated by semantic context, thereby overcoming the static limitations of existing KGC methods.
Figure 3: The architecture of the Flow-Modulated Scoring (FMS) framework. In Semantic Context Learning Module, nodes(square shapes) do not consider their own intrinsic node messages; rather, their role is to facilitate the propagation and aggregation of messages associated with their incident edges, which represent the relation semantics. Conditional Flow Matching Module learns a dynamic flow that modulates the static score to produce the final prediction, unifying static context with dynamic evolution.
3 Methodology
Relational links arise from the interplay of two fundamental principles: the static semantic context of entities and the dynamic evolutionary process connecting them. In this section, we formalize this duality through a probabilistic decomposition of relation prediction (Section 3.1), which in turn motivates our two-component approach. Subsequently, we introduce the core components of the Flow-Modulated Scoring model (Section 3.2) and show how they are unified into a single, cohesive scoring function under a joint training objective (Section 3.2.3).
3.1 Preliminaries
@@
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an instance of a knowledge graph, where $\mathcal{V}$ is the set of entity nodes and $\mathcal{E}$ is the set of edges, with each edge representing a relation $r\in\mathcal{R}$ [26]. Our goal is to predict missing relations in the knowledge graph, i.e., given an entity pair $(h,t)$, infer the relation $r$ between them.
Specifically, we aim to model the distribution of relation types given the entity pair $(h,t)$, denoted as $p(r|h,t)$, to capture the complex semantic context and dynamic nature of relations. According to Bayes’ theorem, this distribution can be expressed as:
$p(r|h,t)\propto p(h,t|r)\cdot p(r),$ (1)
@@
$p(h,t|r)=\frac{1}{2}\left(p(h|r)\cdot p(t|h,r)+p(t|r)\cdot p(h|t,r)\right).$ (2)
@@
Equation (2) provides guidance for our modeling approach. The terms $p(h|r)$ and $p(t|r)$ measure the likelihood of entities $h$ or $t$ given the relation $r$, reflecting the semantic context of the entities. The terms $p(t|h,r)$ and $p(h|t,r)$ capture the dynamic evolution from $h$ to $t$ or from $t$ to $h$ given the relation $r$, reflecting the dynamic nature of relations. The subsequent sections detail how our proposed model instantiates these two factors and integrates them for Knowledge graph completion.
3.2 Flow-Modulated Scoring Model
@@
The overall framework of the Flow-Modulated Scoring model encompasses two core components: Semantic Context Learning and Conditional Flow Matching. As illustrated in Figure 3, the semantic context learning stage is designed to encode context-aware entity embeddings, while the conditional flow matching stage learns a dynamic flow from the head entity embedding to the tail entity embedding. This dynamic flow is then used to modulate a static score for the entity pair, with the ultimate goal of achieving comprehensive relation prediction. Next, we provide a detailed explanation of these components.
3.2.1 Semantic Context Learning